A particle moves along an arc of a circle of radius R according to the law l = a sin ωt, where l is the displacement from the initial position measured along the arc, and a and ω are constants. Assuming R = 1.00 m, a = 0.80 m, and ω = 2.00 rad/s, find: (a) the magnitude of the total acceleration of the particle at the points l = 0 and l = ±a; (b) the minimum value of the total acceleration wmin and the corresponding displacement lm.
1
Expert's answer
2020-10-15T10:36:35-0400

Solution

Acceleration can be written as

"\omega_n=\frac{a^2 \omega^2 \sin\omega t}{R}"

Case first when "l=0" then

"\omega=0, \pi"

Magnitude of acceleration

"\omega_n=\frac{a^2 \omega^2 }{R}=\frac{(0.80)^2 \times2^2 }{1}=2.56 \frac{rad}{sec^2}"

In case of "l=-a" or "+a" then

"\omega =0"

Magnitude of acceleration

"\omega_n=a \omega^2 =(0.80)\times2^2"

"\omega_n=3.20\frac{rad}{sec^2}"